The circumference of a circle and the perimeter of a square are each 20 dm. Which has the greater area, the circle or the square? How much greater?


1
Expert's answer
2021-01-22T01:11:47-0500

The formula for the circumference of a circle is "C = 2\pi R = 20," so "R =10/\pi," then the area is "S= \pi×100/\pi^2 = 100/\pi = 31.83 dm^2."

The side of a square is 20/4 = 5 dm, so its area is "S = 5^2 = 25\,dm^2."

So, the circle has a greater area. It is greater by 31.83 - 25 = 6.83 dm2.